# Emotion-LLaMA: Multimodal Emotion Recognition and Reasoning with Instruction Tuning

### v1

@@ B.3.5 Training Objective and Optimization
B.3.6 Inference Process
During inference, the Emotion-LLaMA model takes as input a multimodal data sample, consisting of visual frames, audio segments, and textual transcriptions. The input data is first processed by the modality-specific encoders, which extract the relevant features and representations. The extracted features are then projected into the unified representation space and concatenated with the appropriate task-specific special tokens. For the emotion recognition task, the model generates a probability distribution over the predefined emotion categories, indicating the likelihood of each emotion being present in the input data. The predicted emotion label is obtained by selecting the category with the highest probability. For the emotion reasoning task, the model generates a natural language explanation that justifies the predicted emotion based on the available multimodal cues. The explanation is generated using a language modeling approach, where the model predicts the most likely sequence of words based on the input data and the learned reasoning patterns. To generate the explanation, we typically use a beam search decoding strategy, which maintains a set of the most promising candidate explanations at each generation step. The beam search algorithm explores multiple hypotheses simultaneously and selects the explanation with the highest cumulative likelihood. The generated explanations provide insights into the model’s reasoning process and highlight the specific cues and evidence that contributed to the predicted emotion. These explanations can be used to interpret the model’s decisions, improve transparency, and facilitate human-machine interaction in emotion-related applications. The online demo is deployed on Hugging Face [^8] https://huggingface.co/spaces/ZebangCheng/Emotion-LLaMA.
@@ C.2.1 Confusion Matrices
Figure 8 presents the confusion matrices for the MER2023 and DFEW datasets, providing insights into the model’s performance across different emotion categories. In both datasets, the categories "angry," "happy," and "sad" are the most common and distinctive emotions encountered in real-world scenarios. These emotions often have clear and pronounced facial expressions, vocal cues, and language patterns, making them relatively easier for the Emotion-LLaMA model to learn and recognize accurately. The confusion matrices show that the model achieves high accuracy for these prevalent emotion categories. The majority of the samples belonging to "angry," "happy," and "sad" are correctly classified, with minimal confusion between them. This indicates that the model has successfully learned the discriminative features and patterns associated with these emotions and can effectively distinguish between them.
However, the confusion matrices also highlight the challenges posed by less frequent emotion categories, such as "disgust" and "fear." Due to the inherent class imbalance present in multimodal emotion recognition datasets, these categories have significantly fewer training samples compared to the more common emotions. The limited availability of training data for "disgust" and "fear" makes it more difficult for the model to learn and generalize the patterns and cues specific to these emotions. As a result, the model may struggle to accurately recognize and classify samples belonging to these categories, leading to higher confusion rates and lower performance compared to the more prevalent emotions. The confusion matrices reveal that a notable proportion of "disgust" and "fear" samples are misclassified as other emotions, such as "angry" or "sad." This suggests that the model may be relying on overlapping or ambiguous features that are shared between these emotions, leading to confusion and misclassification.
@@ C.2.2 Ablation Study of Hyperparameters
C.2.3 Detailed Scoring Cases
@@
Ground Truth In the video, he found out that his friend is pregnant and he was very surprised. He spoke faster, raised his voice, and suddenly sat up from the sofa. He opened his mouth wide, widened his eyes, and raised his eyebrows. Therefore, it can be inferred that his emotional state is surprise.
@@
Video-LLaMA This video showcases a man and a woman in a professional office setting. The woman is speaking to the man, who appears to be a doctor, discussing a case she is handling. They are seated next to a table covered with documents. The woman seems unhappy or stressed while speaking. Throughout the video, the sound of a clock ticking can be heard in the background. The scene is shot in a medium frame with both individuals sitting by the table. The man is wearing a blue shirt and has gray stubble. The woman is in a blue shirt with short hair. Both the table and chairs are made of wood, and the lighting is bright and warm. Overall, the video portrays a professional environment where a doctor is advising a colleague, focusing on their voices and the ambiance of the room.
@@
Emotion-LLaMA In the video, a doctor sits at a desk, displaying facial expressions—a furrowed brow, raised upper eyelids, and pressed lips—that suggest anger or frustration. Along with a tone that implies annoyance or irritation, his words, “How can you talk to Mr. Wang like that?” convey a sense of disapproval and disappointment. This indicates that the person in the video is experiencing anger due to the other person’s misbehavior. Therefore, it is inferred that his emotional state is anger.

### v2

@@ C.2.1 Confusion Matrices
Figure 8 presents the confusion matrices for the MER2023 and DFEW datasets, providing insights into the model’s performance across different emotion categories. In both datasets, the categories “angry,” “happy,” and “sad” are the most common and distinctive emotions encountered in real-world scenarios. These emotions often have clear and pronounced facial expressions, vocal cues, and language patterns, making them relatively easier for the Emotion-LLaMA model to learn and recognize accurately. The confusion matrices show that the model achieves high accuracy for these prevalent emotion categories. The majority of the samples belonging to “angry,” “happy,” and “sad” are correctly classified, with minimal confusion between them. This indicates that the model has successfully learned the discriminative features and patterns associated with these emotions and can effectively distinguish between them.
However, the confusion matrices also highlight the challenges posed by less frequent emotion categories, such as “disgust” and “fear.” Due to the inherent class imbalance present in multimodal emotion recognition datasets, these categories have significantly fewer training samples compared to the more common emotions. The limited availability of training data for “disgust” and “fear” makes it more difficult for the model to learn and generalize the patterns and cues specific to these emotions. As a result, the model may struggle to accurately recognize and classify samples belonging to these categories, leading to higher confusion rates and lower performance compared to the more prevalent emotions. The confusion matrices reveal that a notable proportion of “disgust” and “fear” samples are misclassified as other emotions, such as “angry” or “sad.” This suggests that the model may be relying on overlapping or ambiguous features that are shared between these emotions, leading to confusion and misclassification.
@@ C.2.2 Ablation Study of Hyperparameters
C.2.3 MER2024 Challenge
The MER2024 Challenge [60] has recently garnered significant attention for its focus on integrating multimodalities to identify human emotional states. This initiative aims to address the limitations of existing technologies, which often struggle to meet the demands of practical applications. The challenge seeks to advance research in areas such as multimodal modeling of human affect, modality robustness in affect recognition, low-resource affect recognition, human affect synthesis in multimedia, privacy in affective computing, and applications in health, education, and entertainment. The challenge’s theme has attracted a large number of researchers from around the globe to discuss recent advancements and future directions for robust multimodal emotion recognition, generating many new insights [9, 29, 74, 79].
The MER2024 Challenge has introduced three distinct tracks: (1) MER-SEMI, which provides over 110,000 unlabeled samples for semi-supervised or self-supervised learning; (2) MER-NOISE, which adds noise to test videos to simulate more realistic variations in modality conditions such as background noise and blurry videos, thereby evaluating system robustness; (3) MER-OV, which requires models to extract richer and more subtle emotions in an open-vocabulary manner, aiming to mitigate the intrinsic subjectivity in the emotion annotation process and the inherent ambiguity of MER. Therefore, the MER-NOISE and MER-OV tracks represent the most realistic expressions of emotions and data distributions, presenting both challenges and practical applications.
The MER-NOISE track emphasizes enhancing noise robustness in emotion recognition systems, as noise is ubiquitous in practical settings. Ensuring clear audio streams and high-resolution video frames can be challenging. This track specifically targets two prevalent types of noise: audio additive noise and image blur, encouraging participants to utilize data augmentation techniques and other innovative methods to bolster the resilience of emotion recognition systems [30]. Due to the integration of multimodal information for reasoning, our proposed Emotion-LLaMA significantly reduces the impact of noise, resulting in outstanding performance in the MER-NOISE track with a weighted average F-score (WAF) of 84.52%, surpassing all other teams. Furthermore, we utilized the predictions from Emotion-LLaMA as pseudo-labels to enhance the performance of Conv-Attention [15]. Ultimately, our team leveraged the robust capabilities of Emotion-LLaMA to secure first place in the MER-NOISE track, exceeding the second and third place teams by 1.47% and 1.65%, respectively. Detailed rankings and scores are shown in Figure 10.
Figure 10: Ranking and Scores of Teams in the MER-Noise Track of MER2024.
The MER-OV track introduces the concept of open-vocabulary emotion recognition, addressing the subjectivity and ambiguity inherent in emotion labeling. Traditional datasets often limit label spaces to a few discrete categories, relying on multiple annotators and majority voting, which can overlook valid but non-candidate or minority labels. Participants generate a diverse array of labels, aiming for a more nuanced and accurate representation of emotional states. [33, 98]. As shown in Table 4.4, Emotion-LLaMA demonstrates greater generalization capabilities compared to other multimodal large language models.
C.2.4 Detailed Scoring Cases
@@
Ground Truth In the video, she found out that her friend is pregnant and he was very surprised. She spoke faster, raised her voice, and suddenly sat up from the sofa. She opened her mouth wide, widened her eyes, and raised her eyebrows. Therefore, it can be inferred that her emotional state is surprise.
@@
Video-LLaMA This video showcases a man and a woman in a professional office setting. The woman is speaking to the man, who appears to be a doctor, discussing a case she is handling. They are seated next to a table covered with documents. The woman seems unhappy or stressed while speaking. Throughout the video, the sound of a clock ticking can be heard in the background. The scene is shot in a medium frame with both individuals sitting by the table …
@@
Emotion-LLaMA In the video, a doctor sits at a desk, displaying facial expressions—a furrowed brow, raised upper eyelids—that suggest anger or frustration. Along with a tone that implies annoyance or irritation, his words, “How can you talk to Mr. Wang like that?” convey a sense of disapproval and disappointment. This indicates that the person in the video is experiencing anger due to the other person’s misbehavior. Therefore, it is inferred that his emotional state is anger.
@@ C.3.3 Potential Applications and Impact
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The framework proposed in the paper has been utilized to annotate emotion-related instruction-follow data. Models that are instruction-tuned on this dataset demonstrate robust generalization capabilities. This aligns with the main claims made in the abstract and introduction, accurately reflecting the paper’s contributions and scope.
  5. Guidelines: Sec. 3.1, Table  2
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper addresses the limitations associated with the generation of instruction-follow data, specifically noting that a small subset of the samples exhibits mismatches between emotional cues.
  10. Guidelines: Sec. 3.1, Fiture 4
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper includes formula derivation of the model’s trainable parameters and the principles of model training. The question regarding theoretical results is marked as not applicable (NA) because the paper primarily focuses on how to construct instruction-follow data and how to construct instructions during training.
  15. Guidelines: Sec. 3.1, Sec. 3.3
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper fully discloses all information necessary to reproduce the main experimental results. It includes detailed explanations of the experimental procedures, covering aspects such as hyperparameters and instruction templates.
  20. Guidelines: Sec. 3.3, Sec. 4.2
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: This work provides open access to both the collected instruction-follow data and the trained model parameter files. Detailed descriptions of the necessary steps for reproducing the experiments are included in the paper, ensuring that researchers can faithfully replicate the main experimental results as described in the supplemental material. This facilitates transparency and reproducibility of the research findings.
  25. Guidelines: Sec. 4.2
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper provides a detailed account of the training process, including how training was conducted in phases and all relevant training details. This includes specifics on data splits, hyperparameters, the type of optimizer used, and the instruction templates employed. Such thorough documentation is essential for understanding and validating the results presented in the paper.
  30. Guidelines: Sec. 3.3, Sec. 4.2
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The question regarding the reporting of error bars or statistical significance is marked as not applicable (NA) because the experiments were designed with reproducibility in mind, employing fixed random seeds. Multiple iterations of the experiments consistently yielded results that align with those reported in the paper, ensuring reliability without the explicit use of error bars or detailed statistical tests. This approach ensures that the findings are stable and replicable.
  35. Guidelines: Sec. 4
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper provides detailed information about the computing resources used in the experiments. It specifies the types of compute workers (e.g., GPU models), the memory requirements, and the execution times for each training step.
  40. Guidelines: Sec. 4.2
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This work adheres fully to the NeurIPS Code of Ethics. All datasets employed in the research conform to standards on Copyright and Fair Use, ensuring ethical compliance in terms of data acquisition and utilization. This careful adherence supports the integrity and ethical standards expected in scientific research.
  45. Guidelines: Sec. 4.1
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discusses the potential positive societal impacts of the research, particularly in the fields of human-computer interaction and intelligent education. It explores how the advancements in these areas can enhance user experience and educational outcomes, reflecting on the benefits that the technology can bring to society.
  50. Guidelines: Sec. 1
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: In this work, while generating emotion-related descriptions for unlabeled samples, samples associated with political content and those potentially containing discriminatory material were intentionally filtered out to prevent misuse and negative impacts.
  55. Guidelines: Sec. 3.1
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In this work, all open-source models and datasets utilized adhere to their respective licenses. These resources are appropriately credited within the paper, with explicit mentions of their licenses and terms of use.
  60. Guidelines: Sec. 3.2, Sec. 4.1
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: In this work, the newly generated instruction-follow data is thoroughly documented, including details about its structure, purpose, and the methods used for generation and collection. Comprehensive statistical analyses have been conducted, and all relevant documentation has been made openly available.
  65. Guidelines: Sec. 3.1
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The question regarding crowdsourcing experiments and research with human subjects is marked as not applicable (NA) because the focus of this work is on studying facial expressions using datasets collected and released by other institutions. These datasets comply with all relevant ethical standards. Additionally, this research was conducted under an approved application that adheres to the required regulations for academic studies involving such data.
  70. Guidelines: Sec. 4.1
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The study focuses on the analysis of facial expressions using datasets that were collected and released by other institutions, which had already received IRB approvals. Additionally, this project has obtained the necessary approvals for academic research, ensuring compliance with all relevant regulations.
  75. Guidelines: Sec. 4.1
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.